OpenStudy (anonymous):

put x^2+7x+9=3 into ax^2+bx+c=0? help?

OpenStudy (anonymous):

I got x^2+7x+6=0

OpenStudy (anonymous):

x^2+7x+6=0 so it is equal to ax^2+bx+c=0 therefore a=1,b=7,c=6

OpenStudy (anonymous):

nice! ur right gblover :) keep trying :)